# Discrete homotopy and homology theories for finite posets

### v1

@@ 4.1. Discrete cubical homology
$m^{p}_{t}|\mathcal{K}\{m^{p_{1}}_{t_{1}},\cdots,m^{p_{k}}_{t_{k}}\}|,$
where $a=m_{t}^{p},~{}p_{1}<\dots<p_{k},~{}1\leq t_{j}\leq i_{p_{j}}$, and $|B|=|\mathcal{K}\{m^{p_{1}}_{t_{1}},\cdots,m^{p_{k}}_{t_{k}}\}|$. If
$L(|B|)=L(|A|)=\emptyset,$
@@
$L(b)\not=\emptyset,$
for any $b\in{\rm int}|B|$. In particular, $L(\widehat{B})\not=\emptyset$, where $\widehat{B}$ is the barycenter of $|B|$. Let $C_{1},\cdots,C_{r}$ be all $m$-chains of $M$ containing $\{m^{p_{1}}_{t_{1}},\cdots,m^{p_{k}}_{t_{k}},m^{p}_{t}\}$. Let us take a point $c_{j}\in|\mathcal{K}(C_{j}-m^{p_{k}}_{t_{k}})|$ for each $j$ that is directly below the barycenter of $B$. If $c_{j}\in|\mathcal{K}(f(Q_{n}))|$, then
$|\mathcal{K}(C_{j}-m^{p_{k}}_{t_{k}})|\subseteq|\mathcal{K}(f(Q_{n}))|,$
since $c_{j}$ is by definition in the interior of $|\mathcal{K}(C_{j}-m^{p_{k}}_{t_{k}})|$. This forces
$m^{p_{k}}_{t_{k}}|\mathcal{K}(C_{j}-m^{p_{k}}_{t_{k}})|\subseteq|\mathcal{K}(f(Q_{n}))|,$
$m^{p_{k}}_{t_{k}}$ being a vertex of $f(Q_{n})$, which contradicts the dimension of $|\mathcal{K}(f(Q_{n}))|$. Hence $c_{j}\not\in|\mathcal{K}(f(Q_{n}))|$ and
$|c_{j}A|\cap|\mathcal{K}(f(Q_{n}))|=|A|.$
Let $T_{j}=L_{i}\cup c_{j}A$. Then
$L_{i-1}=\bigcup\limits_{j}T_{j}.$
No points of $|K_{i}|$ are above $|A|$ since this collapse is sunny. It follows that $L_{i}\cap|c_{j}A|=|m^{p}_{t}\partial(c_{j}B)|$, $L_{i}\cup|c_{j}A|=T_{j}$. Each $T_{j}\searrow L_{i}$ is obtained by collapsing $|c_{j}A|$ from $|c_{j}B|$. This finishes the proof that $L_{i-1}\searrow L_{i}$ in this case.
@@
$L(|B|)=\emptyset,~{}L(|A|)\not=\emptyset.$
This says $p=m,~{}p_{k}<m$. We take a point $c^{\prime}_{j}\in|\mathcal{K}(C_{j}-m_{t}^{p})|$ for each $j$ that is directly below $m_{t}^{p}$. The argument for the previous case does apply in this case with $c_{j}$ replaced by $c_{j}^{\prime}$. ∎
Figure 2 shows a poset $M$ and the geometric realization $|\mathcal{K}(M)|$. To illustrate this proof, let $f\in L_{1}(M)$ be represented by $(C,D)$. Then the shaded area in the figure on the right is $L_{0}$.
Figure 2.
Observe that an $(m,q)$-ball pair $(B^{m},B^{q})$ is unknotted by the argument in the proof of [12, Theorem 9], provided that $B^{m}\searrow B^{q}$. If $\partial|\mathcal{K}(f(Q_{n}))|\subseteq\partial|\mathcal{K}(M)|$ and ${\rm int}|\mathcal{K}(f(Q_{n}))|\subseteq{\rm int}|\mathcal{K}(M)|$, $(|\mathcal{K}(M)|,|\mathcal{K}(f(Q_{n}))|)$ is a ball pair. Thus by Lemma 4.13, the following result holds.
@@
Let $M$ be a contractible poset with $|\mathcal{K}(M)|$ being an $m$-manifold and $f\in L_{n}(M)$. If ${\rm dim}f(Q_{n})<m$ and $(|\mathcal{K}(M)|,|\mathcal{K}(f(Q_{n}))|)$ is a ball pair, then
$(|\mathcal{K}(M)|,|\mathcal{K}(f(Q_{n}))|)$
@@
Let $M$ be a contractible poset with $|\mathcal{K}(M)|$ being an $m$-manifold and $f_{1}\in L_{n}(M)$ with $|\mathcal{K}(f_{1}(Q_{n}))|\subseteq{\rm int}|\mathcal{K}(M)|$. Assume that ${\rm dim}f_{1}(Q_{n})<m$ and $(|\mathcal{K}(M)|,|\mathcal{K}(f(Q_{n}))|)$ is a ball pair. Let $c=dim|\mathcal{K}(M-f_{1}(Q_{n}))|-1$. If $n\leq m-1$ and $p\leq c$, then for any $f_{2}\in L_{p}(M-f_{1}(Q_{n}))$ with $|\mathcal{K}(f_{2}(Q_{p}))|\subseteq{\rm int}|\mathcal{K}(M-f_{1}(Q_{n}))|$ and $x\in f_{1}(Q_{n})$, the $p$th Betti number of
$M-f_{1}(Q_{n})-f_{2}(Q_{p})$
@@
$\psi_{*}:{H}_{p}^{\rm Cube}(M-f_{1}(Q_{n})-f_{2}(Q_{p}))\rightarrow{H}_{p}^{\rm Simpl}(\mathcal{K}(M-f_{1}(Q_{n})))$
is surjective. Because $(|\mathcal{K}(M)|,|\mathcal{K}(f_{1}(Q_{n}))|)$ in an unknotted ball pair,
${H}_{q}^{\rm Sing}(|\mathcal{K}(M)|-|\mathcal{K}(f_{1}(Q_{n}))|)=\begin{cases}\mathbb{Z},~{}&\text{if}~{}q=m-2,~{}0<n<m-1,~{}\text{or}\\ &~{}q=m-1,~{}n=0,~{}\text{or}~{}q=0,\\ 0,~{}&\text{otherwise}.\end{cases}$
@@
Example 4.16.
Let $M$ be the following poset, which is a finite model of the $3$-ball.
${i}$${f\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${g\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${h\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${c\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${d\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${e\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${a\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${b\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$.
Let $f_{1}\in L_{0}(M)$ be the constant map $Q_{0}\rightarrow i$, and let $f_{2}\in L_{1}(M-f_{1}(Q_{0}))$ be the represented $(b,c)$. Then $M-f_{1}(Q_{0})-f_{2}(Q_{1})$ is a finite model of the circle, being not contractible.
@@
Let $M=|K|$ be a triangulated $m$-manifold with $K$ simplicially collapsible and $f_{1}\in L_{n}(\chi(K))$ with $|\mathcal{K}(f_{1}(Q_{n}))|\subseteq{\rm int}M$. Assume also that ${\rm dim}f_{1}(Q_{n})<m$ and $(M,|\mathcal{K}(f(Q_{n}))|)$ is a ball pair. Let $c=dim|\mathcal{K}(\chi(K)-f_{1}(Q_{n}))|-1$ If $n\leq m-1$, then for any $f_{2}\in L_{p}(\chi(K)-f_{1}(Q_{n}))$ with $|\mathcal{K}(f_{2}(Q_{p}))|\subseteq{\rm int}|\mathcal{K}(\chi(K)-f_{1}(Q_{n}))|$,
$M-|\mathcal{K}(f_{1}(Q_{n}))|-|\mathcal{K}(f_{2}(Q_{n}))|$
in not contractible, if $10<p=m-2\leq c,~{}0<n<m-$ 1, or $00<p=m-1\leq c,~{}n=$ 0.
@@ 5. Discrete Hurewicz map
The aim of this section is to construct an analogue of the Hurewicz homomorphism connecting discrete homotopy theory with discrete cubical homology theory. The discrete Hurewicz homomorphism that we define is similar to the one in [8]. As stated in Theorem 5.5, this discrete Hurewicz homomorphism indeed coincides with the classical one,
Given a finite poset $X$ with basepoint $x_{0}$. Let $f:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ be a map of radius $r_{f}$. Define $H:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\times I_{r}\rightarrow(X,x_{0})$ by $H(x,i)=f(x_{1}-i,\cdots,x_{n}-i)$ for $(x,i)\in\mathbb{Z}^{n}\times I_{r}$, where $r>r_{f}$. Then $H$ is a basepoint-preserving-homotopy from $f$ to a map $H_{r}$. Take $T_{r}f:(I_{2r}^{n},\partial I_{2r}^{n})\rightarrow(X,x_{0})$ to be the restriction of $H_{r}$ to $I^{n}_{2r}$.
@@
For $1\leq i\leq n$, let $\Gamma_{i}=\{x\in I_{2r-1}^{n}~{}|~{}x_{i}=0\}$. One has
$\begin{split}\partial_{n}^{\rm Cube}\varphi(T_{r}f)&=\sum_{i=1}^{n}(-1)^{i}\sum_{x\in\Gamma_{i}}\sum_{r=0}^{2r-1}\left(f_{i}^{-}(T_{r}f)^{x+re_{i}}-f_{i}^{+}(T_{r}f)^{x+re_{i}}\right)\\ &=\sum_{i=1}^{n}(-1)^{i}\sum_{x\in\Gamma_{i}}\left(f_{i}^{-}(T_{r}f)^{x}-f_{i}^{+}(T_{r}f)^{x+(2r-1)e_{i}}\right),\end{split}$
where $e_{i}\in\mathbb{Z}^{n}$ ia a vector with $1$ in the $i$-th position and 00 everywhere else. Observe that for each $i$ and $x\in\Gamma_{i}$, $f_{i}^{-}(T_{r}f)^{x}-f_{i}^{+}(T_{r}f)^{x+(2r-1)e_{i}}$ is 00 since $T_{r}f$ is constant on $\partial I_{2r}^{n}$. This proves $(2)$.
@@
$\begin{split}I^{n}_{2r}\times I_{p}&\rightarrow X,\\ (x,t)&\mapsto H(x-r,t),\end{split}$
$r$ being taken the maximum integer of radii of $H_{i},~{}0\leq i\leq p$. Let $\Gamma_{i}=\{x\in I^{n}_{2r-1}\times I_{p-1}~{}|~{}x_{i}=0\}$ and let $e_{i}\in\mathbb{Z}^{n+1}$ be the vector with $1$ in the $i$-th position and 00 everywhere else. We have
$\begin{split}\partial_{n+1}^{\rm Cube}\varphi(TH)(y)&=\sum_{i=1}^{n}(-1)^{i}\sum_{x\in\Gamma_{i}}\sum_{r=0}^{2r-1}\left(f_{i}^{-}(TH)^{x+re_{i}}(y)-f_{i}^{+}(TH)^{x+re_{i}}(y)\right)\\ +&(-1)^{n+1}\sum_{x\in\Gamma_{n+1}}\sum_{r=0}^{p-1}\left(f_{n+1}^{-}(TH)^{x+re_{n+1}}(y)-f_{n+1}^{+}(TH)^{x+re_{n+1}}(y)\right)\\ &=(-1)^{n+1}\sum_{x\in\Gamma_{n+1}}\left(f_{n+1}^{-}(TH)^{x}(y)-f_{n+1}^{+}(TH)^{x+(p-1)e_{n+1}}(y)\right)\\ &=(-1)^{n+1}\sum_{x\in\Gamma_{n+1}}(TH)(x_{1}+y_{1},\cdots,x_{n}+y_{n},0)\\ -&(-1)^{n+1}\sum_{x\in\Gamma_{n+1}}(TH)(x_{1}+y_{1},\cdots,x_{n}+y_{n},p)\\ &=(-1)^{n+1}\left(\varphi({(TH)}_{0})-\varphi({(TH)}_{p})\right)(y)\\ &=(-1)^{n+1}\left(\varphi(T_{r_{f}}f)-\varphi(T_{r_{g}}g)\right)(y),\end{split}$
@@
There is a discrete Hurewicz theorem in dimension $1$. The same strategy in the proof [4, Theorem 4.1] applies in this case.
@@
$\begin{split}\eta_{*}\psi_{*}h^{D}([f])&=\eta_{*}\psi[\sum_{x\in I^{n}_{2r-1}}(T_{r}f)^{x}]\\ &=\eta_{*}\left[\sum_{x\in I^{n}_{2r-1}}\sum_{\tau\in S_{n}}{\rm sign}(\tau)\mathcal{K}\left(\{(T_{r}f)^{x}(p_{\tau}(i))~{}|~{}0\leq i\leq n\}\right)\right]\\ &=\eta_{*}\left[\sum_{x\in I^{n}_{2r-1}}\sum_{\tau\in S_{n}}{\rm sign}(\tau)\mathcal{K}\left(\{T_{r}f(x+p_{\tau}(i))~{}|~{}0\leq i\leq n\}\right)\right]\\ &=\eta_{*}\mathcal{K}(T_{r}f)_{*}\left[\sum_{x\in I^{n}_{2r-1}}\sum_{\tau\in S_{n}}{\rm sign}(\tau)\mathcal{K}\left(\{x+p_{\tau}(i)~{}|~{}0\leq i\leq n\}\right)\right],\end{split}$
where $\left[\sum\limits_{x\in I^{n}_{2r-1}}\sum\limits_{\tau\in S_{n}}{\rm sign}(\tau)\mathcal{K}\left(\{x+p_{\tau}(i)~{}|~{}0\leq i\leq n\}\right)\right]$ is a generator of
$H^{\rm Simpl}_{n}(\mathcal{K}(I^{n}_{2r}),\mathcal{K}(\partial I^{n}_{2r})),$
@@
$P(x)=x+r.$
Then we have $f(x)=T_{r}f(Px)$ for every $x\in\mathbb{Z}_{\leq r}^{n}$. This says when restricted to $\mathbb{Z}_{\leq r}^{n}$,
$|\mathcal{K}(f)|=|\mathcal{K}(T_{r}f)\circ\mathcal{K}(P)|.$
Denote by $\beta={\lambda_{r}}_{*}|\mathcal{K}(P|\mathbb{Z}_{\leq 2r})|^{-1}_{*}\eta_{*}\alpha$ the fixed generator of $H^{\rm Sing}_{n}(L^{n},\partial L^{n})$. Then
$h\theta([f])$ $=(|\mathcal{K}(f)|\circ\lambda_{r}^{-1})_{*}\eta_{*}(\beta)$
$=|\mathcal{K}(T_{r}f)\circ\mathcal{K}(P|\mathbb{Z}_{\leq 2r})|_{*}\circ{(\lambda_{r}^{-1})}_{*}({(\lambda_{r})}_{*}|\mathcal{K}(P|\mathbb{Z}_{\leq 2r})|^{-1}_{*}\eta_{*}\alpha)$

### v2

@@ 4.1. Discrete cubical homology
$x^{p}_{t}|\mathcal{K}\{x^{p_{1}}_{t_{1}},\cdots,x^{p_{k}}_{t_{k}}\}|,$
where $a=x_{t}^{p},~p_{1}<\dots<p_{k},~1\leq t_{j}\leq i_{p_{j}}$, and $B=|\mathcal{K}\{x^{p_{1}}_{t_{1}},\cdots,x^{p_{k}}_{t_{k}}\}|$. If
$E(B)=E(A)=\emptyset,$
@@
$E(b)\not$$\emptyset$,
[…]
$|\mathcal{K}(C_{j}-x^{p_{k}}_{t_{k}})|\subseteq P_{0},$
since $c_{j}$ is by definition in the interior of $|\mathcal{K}(C_{j}-x^{p_{k}}_{t_{k}})|$. This forces
$x^{p_{k}}_{t_{k}}|\mathcal{K}(C_{j}-x^{p_{k}}_{t_{k}})|\subseteq P_{0},$
$x^{p_{k}}_{t_{k}}$ being a vertex of $\sigma(Q_{n})$, which contradicts the dimension of $P_{0}$. Hence $c_{j}\not\in P_{0}$ and
$c_{j}A\cap P_{0}=A.$
Let $T_{j}=W_{i}\cup c_{j}A$. Then
$W_{i-1}=\bigcup\limits_{j}T_{j}.$
No points of $P_{i}$ are above $A$ since this collapse is sunny. It follows that $W_{i}\cap c_{j}A=|x^{p}_{t}\partial(c_{j}B)|$, $W_{i}\cup c_{j}A=T_{j}$. Each $T_{j}\searrow W_{i}$ is obtained by collapsing $c_{j}A$ from $c_{j}B$. This finishes the proof that $W_{i-1}\searrow W_{i}$ in this case.
@@
$E(B)=\emptyset,~E(A)\not$$\emptyset$.
This says $p=m,~p_{k}<m$. We take a point $c^{\prime}_{j}\in|\mathcal{K}(C_{j}-x_{t}^{p})|$ for each $j$ that is directly below $x_{t}^{p}$. The argument for the previous case does apply in this case with $c_{j}$ replaced by $c_{j}^{\prime}$. ∎
Figure 5 shows a poset $X$ and the geometric realization $|\mathcal{K}(X)|$. To illustrate this proof, let $\sigma\in L_{1}(X)$ be represented by $(C,D)$. Then the shaded area in the figure on the right is $W_{0}$.
Figure 5. A poset $X$ and the geometric realization $|\mathcal{K}(X)|$.
Observe that an $(m,q)$-ball pair $(B^{m},B^{q})$ is unknotted by the argument in the proof of [17, Theorem 9], provided that $B^{m}\searrow B^{q}$. If $\partial|\mathcal{K}(\sigma(Q_{n}))|\subseteq\partial|\mathcal{K}(X)|$ and ${\rm int}|\mathcal{K}(\sigma(Q_{n}))|\subseteq{\rm int}|\mathcal{K}(X)|$,
$(|\mathcal{K}(X)|,|\mathcal{K}(\sigma(Q_{n}))|)$
is a ball pair. Thus by Lemma 4.16, the following result holds.
@@
Let $X$ be a contractible poset with $|\mathcal{K}(X)|$ being an $m$-manifold and $\sigma\in L_{n}(X)$. If ${\rm dim}\sigma(Q_{n})<m$ and $(|\mathcal{K}(X)|,|\mathcal{K}(\sigma(Q_{n}))|)$ is a ball pair, then
$(|\mathcal{K}(X)|,|\mathcal{K}(\sigma(Q_{n}))|)$
@@
Let $X$ be a contractible poset with $|\mathcal{K}(X)|$ being an $m$-manifold and $\sigma_{1}\in L_{n}(X)$ with $|\mathcal{K}(\sigma_{1}(Q_{n}))|\subseteq{\rm int}|\mathcal{K}(X)|$. Assume that ${\rm dim}\sigma_{1}(Q_{n})<m$ and $(|\mathcal{K}(X)|,|\mathcal{K}(\sigma_{1}(Q_{n}))|)$ is a ball pair. Let
$c=dim|\mathcal{K}(X-\sigma_{1}(Q_{n}))|-1.$
If $n\leq m-1$ and $p\leq c$, then for any $\sigma_{2}\in L_{p}(X-\sigma_{1}(Q_{n}))$ with $|\mathcal{K}(\sigma_{2}(Q_{p}))|\subseteq{\rm int}|\mathcal{K}(X-\sigma_{1}(Q_{n}))|$ and $x\in\sigma_{1}(Q_{n})$, the $p$th Betti number of
$X-\sigma_{1}(Q_{n})-\sigma_{2}(Q_{p})$
@@
$\psi_{*}:{H}_{p}^{\rm Cube}(X-\sigma_{1}(Q_{n})-\sigma_{2}(Q_{p}))\rightarrow{H}_{p}^{\rm Simpl}(\mathcal{K}(X-\sigma_{1}(Q_{n})))$
is surjective. From Proposition 4.17, we know that $(|\mathcal{K}(X)|,|\mathcal{K}(\sigma_{1}(Q_{n}))|)$ in an unknotted ball pair. Then
${H}_{q}^{\rm Sing}(|\mathcal{K}(X)|-|\mathcal{K}(\sigma_{1}(Q_{n}))|)=\begin{cases}\mathbb{Z},~&\text{if}~q=m-2,~0<n<m-1,~\text{or}\\ &~q=m-1,~n=0,~\text{or}~q=0,\\ 0,~&\text{otherwise}.\end{cases}$
@@
Let $M=|K|$ be a triangulated $m$-manifold with $K$ simplicially collapsible and $\sigma_{1}\in L_{n}(\chi(K))$ with $|\mathcal{K}(\sigma_{1}(Q_{n}))|\subseteq{\rm int}M$. Assume also that ${\rm dim}\sigma_{1}(Q_{n})<m$ and $(M,|\mathcal{K}(\sigma_{1}(Q_{n}))|)$ is a ball pair. Let
$c=dim|\mathcal{K}(\chi(K)-\sigma_{1}(Q_{n}))|-1.$
If $n\leq m-1$, then for any $\sigma_{2}\in L_{p}(\chi(K)-\sigma_{1}(Q_{n}))$ with $|\mathcal{K}(\sigma_{2}(Q_{p}))|\subseteq{\rm int}|\mathcal{K}(\chi(K)-\sigma_{1}(Q_{n}))|$,
$M-|\mathcal{K}(\sigma_{1}(Q_{n}))|-|\mathcal{K}(\sigma_{2}(Q_{n}))|$
in not contractible, if $0<p=m-2\leq c,~0<n<m-1$, or $0<p=m-1\leq c,~n=0$.
The next example will help better understand the content of Theorem 4.18.
Example 4.20.
Let $X$ be the following poset, which is a finite model of the $3$-ball.
[…]
Let $\sigma_{1}\in L_{0}(X)$ be the constant map $Q_{0}\rightarrow i$, and let $\sigma_{2}\in L_{1}(X-\sigma_{1}(Q_{0}))$ be the represented $(b,c)$. Then $X-\sigma_{1}(Q_{0})-\sigma_{2}(Q_{1})$ is a finite model of the circle, being not contractible.
@@ 5. Discrete Hurewicz map
The aim of this section is to construct an analogue of the Hurewicz homomorphism connecting discrete homotopy theory with discrete cubical homology theory. The discrete Hurewicz homomorphism to be defined is similar to the one in [11]. This discrete Hurewicz homomorphism indeed coincides with the classical one. This is stated in Theorem 5.5.
Given a finite poset $X$ with basepoint $x_{0}$, let $f:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ be a map of radius $r$. Define $f^{\prime}:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ by
$f^{\prime}(x)=f(x_{1}-2r,\cdots,x_{n}-2r).$
Take $T_{r}f:(I_{4r}^{n},\partial I_{4r}^{n})\rightarrow(X,x_{0})$ to be the restriction of $f^{\prime}$ to $I^{n}_{4r}$.
@@
For $1\leq i\leq n$, let $\Gamma_{i}=\{x\in I_{4r-1}^{n}~|~x_{i}=0\}$. One has
$\begin{split}\partial_{n}^{\rm Cube}\varphi(T_{r}f)&=\sum_{i=1}^{n}(-1)^{i}\sum_{x\in\Gamma_{i}}\sum_{r=0}^{4r-1}\left(f_{i}^{-}(T_{r}f)^{x+re_{i}}-f_{i}^{+}(T_{r}f)^{x+re_{i}}\right)\\ &=\sum_{i=1}^{n}(-1)^{i}\sum_{x\in\Gamma_{i}}\left(f_{i}^{-}(T_{r}f)^{x}-f_{i}^{+}(T_{r}f)^{x+(4r-1)e_{i}}\right),\end{split}$
where $e_{i}\in\mathbb{Z}^{n}$ is a vector with $1$ in the $i$-th position and 0 everywhere else. Observe that for each $i$ and $x\in\Gamma_{i}$, $f_{i}^{-}(T_{r}f)^{x}-f_{i}^{+}(T_{r}f)^{x+(4r-1)e_{i}}$ is 0 since $T_{r}f$ is constant on $\partial I_{4r}^{n}$. This proves $(2)$.
@@
$\begin{split}I^{n}_{4r}\times I_{p}&\rightarrow X,\\ (x,t)&\mapsto H(x-2r,t),\end{split}$
$r$ being taken the maximum integer of radii of $H_{i},~0\leq i\leq p$. Let
$\Gamma_{i}=\{x\in I^{n}_{4r-1}\times I_{p-1}~|~x_{i}=0\}$
and let $e_{i}\in\mathbb{Z}^{n+1}$ be the vector with $1$ in the $i$-th position and 0 everywhere else. We have
$\begin{split}\partial_{n+1}^{\rm Cube}\varphi(TH)(y)&=\sum_{i=1}^{n}(-1)^{i}\sum_{x\in\Gamma_{i}}\sum_{r=0}^{4r-1}\left(f_{i}^{-}(TH)^{x+re_{i}}(y)-f_{i}^{+}(TH)^{x+re_{i}}(y)\right)\\ +&(-1)^{n+1}\sum_{x\in\Gamma_{n+1}}\sum_{r=0}^{p-1}\left(f_{n+1}^{-}(TH)^{x+re_{n+1}}(y)-f_{n+1}^{+}(TH)^{x+re_{n+1}}(y)\right)\\ &=(-1)^{n+1}\sum_{x\in\Gamma_{n+1}}\left(f_{n+1}^{-}(TH)^{x}(y)-f_{n+1}^{+}(TH)^{x+(p-1)e_{n+1}}(y)\right)\\ &=(-1)^{n+1}\sum_{x\in\Gamma_{n+1}}(TH)(x_{1}+y_{1},\cdots,x_{n}+y_{n},0)\\ -&(-1)^{n+1}\sum_{x\in\Gamma_{n+1}}(TH)(x_{1}+y_{1},\cdots,x_{n}+y_{n},p)\\ &=(-1)^{n+1}\left(\varphi({(TH)}_{0})-\varphi({(TH)}_{p})\right)(y)\\ &=(-1)^{n+1}\left(\varphi(T_{r_{f}}f)-\varphi(T_{r_{g}}g)\right)(y),\end{split}$
@@
From Lemma 5.2, we know that $h^{D}$ is well-defined. The next proposition will show that indeed $h^{D}$ preserves group structure.
@@
There is a discrete Hurewicz theorem in dimension $1$. The same strategy in the proof [3, Theorem 4.1] applies in this case.
@@
$\begin{split}\eta_{*}\psi_{*}h^{D}([f])&=\eta_{*}\psi[\sum_{x\in I^{n}_{4r-1}}(T_{r}f)^{x}]\\ &=\eta_{*}\left[\sum_{x\in I^{n}_{4r-1}}\sum_{\tau\in S_{n}}{\rm sign}(\tau)\mathcal{K}\left(\{(T_{r}f)^{x}(p_{\tau}(i))~|~0\leq i\leq n\}\right)\right]\\ &=\eta_{*}\left[\sum_{x\in I^{n}_{4r-1}}\sum_{\tau\in S_{n}}{\rm sign}(\tau)\mathcal{K}\left(\{T_{r}f(x+p_{\tau}(i))~|~0\leq i\leq n\}\right)\right]\\ &=\eta_{*}\mathcal{K}(T_{r}f)_{*}\left[\sum_{x\in I^{n}_{4r-1}}\sum_{\tau\in S_{n}}{\rm sign}(\tau)\mathcal{K}\left(\{x+p_{\tau}(i)~|~0\leq i\leq n\}\right)\right],\end{split}$
where $\left[\sum\limits_{x\in I^{n}_{4r-1}}\sum\limits_{\tau\in S_{n}}{\rm sign}(\tau)\mathcal{K}\left(\{x+p_{\tau}(i)~|~0\leq i\leq n\}\right)\right]$ is a generator of
$H^{\rm Simpl}_{n}(\mathcal{K}(I^{n}_{4r}),\mathcal{K}(\partial I^{n}_{4r})),$
@@
$P(x)=x+2r.$
Then we have $f(x)=T_{r}f(Px)$ for every $x\in\mathbb{Z}_{[-2r,2r]}^{n}$. It follows that
$|\mathcal{K}(f)|=|\mathcal{K}(T_{r}f)\circ\mathcal{K}(P)|$
when restricted to $\mathbb{Z}_{[-2r,2r]}^{n}$. Denote by $\beta={\lambda_{2r}}_{*}|\mathcal{K}(P)|^{-1}_{*}\eta_{*}\alpha$ the fixed generator of $H^{\rm Sing}_{n}(L^{n},\partial L^{n})$. Then
$h\theta([f])$ $=(|\mathcal{K}(f)|\circ\lambda_{2r}^{-1})_{*}\eta_{*}(\beta)$
$=|\mathcal{K}(T_{r}f)\circ\mathcal{K}(P)|_{*}\circ{(\lambda_{2r}^{-1})}_{*}({(\lambda_{2r})}_{*}|\mathcal{K}(P)|^{-1}_{*}\eta_{*}\alpha)$